OpenStudy (migitmack):

Jimmy's Chocolate shop sells two different fudge bars. The dark chocolate fudge bars sell for $5 per pound and the milk chocolate fudge bars sell for $4 per pound. You purchased a combination of these bars for your family and friends and your bill came to $150. This scenario can be modeled by the equation 5d + 4m = 150. Which of the following is not a possible combination of dark chocolate fudge bars and milk chocolate fudge bars? 6 dark bars and 30 milk bars 10 dark bars and 25 milk bars 12 dark bars and 23 milk bars 14 dark bars and 20 milk bars

OpenStudy (anonymous):

c

OpenStudy (anonymous):

3) solution: d=6 m=30 5d + 4m = 30 + 120 = 150 (possible) d=10 m=25 5d + 4m = 50 + 100 = 150 (possible) d=12 m=23 5d + 4m = 60 + 92 = 152 (not possible) d=14 m=20 5d + 4m = 70 + 80 = 150 (possible)